# Horizon Brightened Acceleration Radiation from Massive Vector Fields

### v1

@@ VII.2 Area-entropy relation with Proca modifications
We now relate the entropy flux carried by the escaping Proca quanta to the change of the black-hole area. Summing (100) over all cavity-admitted channels gives the total entropy rate
$\frac{dS_{p}}{dt}=4\pi k_{B}\sum_{\nu,\ell}\nu\,\dot{\bar{n}}_{\nu}(\ell).$ (101)
@@
$\dot{E}_{V}=\hbar\,\frac{c}{r_{g}}\sum_{\nu,\ell}\nu\,\dot{\bar{n}}_{\nu}(\ell)=\hbar\,\frac{c^{3}}{2GM}\sum_{\nu,\ell}\nu\,\dot{\bar{n}}_{\nu}(\ell).$ (102)
@@
$\dot{A}=\frac{32\pi G^{2}M}{c^{4}}\,\dot{M}.$ (103)
Combining (101)-(103) eliminates the mode sum and yields
$\frac{dS_{p}}{dt}=-\,\frac{k_{B}c^{3}}{4\hbar G}\,\dot{A}.$ (104)
@@
$\frac{dS_{p}}{dt}\;=\;\frac{k_{B}c^{3}}{4\hbar G}\,\dot{A}_{V}\;.$ (106)
Equation (106) is universal: the proportionality constant depends only on $(G,\hbar,c,\,k_{B})$ and the near-horizon mapping that fixed detailed balance, while all Proca-specific physics (mass threshold $\nu\geq m_{V}$, polarization sums, axial/polar greybody factors, and detector realization) enters solely through $\dot{A}_{V}$ via the net flux $\sum_{\nu,\ell}\nu\,\dot{\bar{n}}_{\nu}(\ell)$ in (102). In the massless limit $m_{V}\to 0$ the same relation holds, with $\dot{A}_{V}$ computed from the corresponding threshold-free spectrum.
@@ VII.3 Spectral signatures and limits
We distill the qualitative and asymptotic features of the Proca spectrum that enter the entropy flux through $\dot{\bar{n}}_{\nu}$ and, hence, $\dot{S}_{p}$ in (100)-(106). The thermal denominator is general; all structure resides in the smooth prefactor and the kinematic threshold $\nu\geq m_{V}$.
@@
The massless limit $m_{V}\to 0$ removes the threshold and sends $k_{\infty}\to\nu$, returning the scalar-like normalization $1/\nu$ (up to polarization weights). The dipole model passes smoothly to the standard electromagnetic response. The current model, while formulated naturally for Proca, still yields a smooth limit at the level of observable probabilities and rates. Finally, all results here rest on the large-$\omega$ expansion that produced the universal near-horizon integral; corrections scale as $\mathcal{O}(\nu/\omega)$, reshaping only the smooth prefactor and leaving intact the Planckian denominator and detailed balance. Consequently, the entropy flux $\dot{S}_{p}=4\pi k_{B}\sum_{\nu,\ell}\nu\dot{\bar{n}}_{\nu}(\ell)$ inherits the same signatures: delayed onset at $\nu\approx m_{V}$, gentle turn-on dominated by polar $\ell=0$, and a universal thermal tail at higher frequencies.
These signatures (sharp threshold, channel-dependent turn-on, polarization-weighted prefactors, and a universal thermal tail) are the distinctive Proca imprints in the HBAR entropy law (106).
@@ VIII.1 Large-$\omega$ expansion and validity
$\mathcal{A}_{\bullet}(\nu,\ell,m,\lambda)\;=\;\int d\tau\;\mathcal{G}_{\bullet}(\tau,\nu,\ell,m,\lambda)\,e^{\,i\Phi(\tau,\nu)}.$ (107)
@@
$\Phi(\tau,\nu)=\omega\tau+\nu\,\left[t(\tau)-r_{*}(\tau)\right]+\Phi_{0}(\nu),$ (108)
@@
$\mathcal{I}(\nu,\omega)\;\equiv\;\int_{0}^{\infty}dy\,y^{-2i\nu}\,e^{-i\omega y}\;=\;(i\omega)^{-(1-2i\nu)}\Gamma(1-2i\nu),$ (109)
@@
$|\mathcal{I}(\nu,\omega)|^{2}=\frac{4\pi\,\nu}{\omega^{2}}\,\frac{1}{e^{4\pi\nu}-1}.$ (110)
The large-$\omega$ control arises because the integrand oscillates rapidly on a scale $y\sim\omega^{-1}$. Provided the envelope is smooth on this scale, we can replace $\mathcal{G}_{\bullet}(\tau_{h}-y,\nu,\ell,m,\lambda)$ by its horizon value. A Taylor expansion yields
$\mathcal{G}_{\bullet}(\tau_{h}-y,\nu,{\cdot})=\mathcal{G}_{\bullet h}(\nu,{\cdot})-\mathcal{G}^{\prime}_{\bullet h}(\nu,{\cdot})\,y+\frac{1}{2}\mathcal{G}^{\prime\prime}_{\bullet h}(\nu,{\cdot})\,y^{2}+\cdots,$ (111)
and term-wise integration gives a controlled asymptotic series,
$\mathcal{A}_{\bullet}=\mathcal{G}_{\bullet h}\,\mathcal{I}-\mathcal{G}^{\prime}_{\bullet h}\,\partial_{(i\omega)}\mathcal{I}+\frac{1}{2}\mathcal{G}^{\prime\prime}_{\bullet h}\,\partial_{(i\omega)}^{2}\mathcal{I}+\cdots.$ (112)
Using $\partial_{(i\omega)}^{n}\mathcal{I}\sim\omega^{-1-n+2i\nu}\Gamma(1+n-2i\nu)$, one finds a relative expansion
$\frac{\mathcal{A}_{\bullet}}{\mathcal{G}_{\bullet h}\,\mathcal{I}}=1+\mathcal{O}\!\left(\frac{\nu}{\omega}\right),\qquad\frac{P_{\rm exc}^{\bullet}}{|\mathcal{G}_{\bullet h}|^{2}\,|\mathcal{I}|^{2}}=1+\mathcal{O}\!\left(\frac{\nu}{\omega}\right).$ (113)
Thus, the Planck factor in $|\mathcal{I}|^{2}$ is stable; corrections only reshape the prefactor at order $\nu/\omega$.
@@
$\omega\,\delta\tau_{\rm cav}\gg 1\quad\Rightarrow\quad\text{single-pass, near-horizon dominance.}$ (114)
@@ VIII.2 Gauge and constraints
The Proca theory is massive and therefore has no gauge symmetry. The covariant Lorenz condition $\nabla_{\mu}A^{\mu}=0$ follows dynamically from the equations of motion, leaving three propagating degrees of freedom. This structure percolates cleanly through mode expansion and detector couplings. Asymptotically, the three polarizations satisfy $k\!\cdot\!\varepsilon^{(\lambda)}=0$, and the projector is tong2014quantum
$\sum_{\lambda=1}^{3}\varepsilon^{(\lambda)}_{\mu}\varepsilon^{(\lambda)}_{\nu}=-\eta_{\mu\nu}+\frac{k_{\mu}k_{\nu}}{m_{V}^{2}}.$ (115)
Near the horizon, the axial/polar split maps to three independent master variables with plane-wave behaviour $e^{-i\nu(t-r_{*})}$; the constraint does not modify the universal phase.
@@
$u\!\cdot\!u+\frac{(u\!\cdot\!k)^{2}}{m_{V}^{2}}=\,-1+\frac{\Omega^{2}}{m_{V}^{2}},$ (116)
@@
$E_{\mu}\equiv F_{\mu\nu}u^{\nu}=i\left[k_{\mu}(A\!\cdot\!u)-(k\!\cdot\!u)\,A_{\mu}\right]$ (117)
automatically satisfies $u^{\mu}E_{\mu}=0$. Both pieces contribute at the amplitude level; after summing polarizations the weight is built from contractions with the projector in (115). Orientation averaging of $d^{\mu}$ in the atom’s rest frame produces a smooth, $\mathcal{O}(1)$ function of $\nu$.
Finally, for the massless limit, as $m_{V}\to 0$, the projector (115) becomes singular, but all observables remain finite: the combination of unit-flux normalization, $\Xi_{\ell}(\nu)$, and the longitudinal/transverse balance yields a smooth limit in probabilities and rates. Physically, the dipole model tends to the familiar electromagnetic response; the current model, while not gauge-invariant in the $m_{V}=0$ theory, approaches the same observables because the longitudinal contribution decouples as $\nu$ grows and $\Xi_{\ell}\to 1$.
@@ VIII.3 Comparison with Unruh/DeWitt logic
Our calculation is the black-hole (Killing/Rindler) analogue of the Unruh-DeWitt (UDW) detector response. Three points make the connection precise. First, in the UDW setting, a uniformly accelerated detector with proper acceleration $a$ sees a thermal spectrum with temperature $T=a/(2\pi)$, traced to the KMS periodicity of the Wightman function in imaginary proper time. Here, along the free-fall trajectory, the near-horizon mapping
$t(\tau)-r_{*}(\tau)=-\,2\ln(\tau_{h}-\tau)+C_{h}+\cdots$ (118)
acts as a local Rindler time. The logarithm produces a branch cut whose discontinuity across imaginary $2\pi$ shifts enforces Unruh:1976db; Hartle:1976tp
$\frac{P_{\rm exc}(\nu)}{P_{\rm abs}(\nu)}=e^{-4\pi\nu},$ (119)
exactly the KMS ratio with surface gravity $\kappa=\frac{1}{2}$ in our dimensionless units. This universality explains why spin, mass, and detector realization cannot change the thermal denominator. Second, in the Boulware vacuum, the rotating term $\propto a_{\nu\ell m\lambda}\sigma_{eg}$ vanishes on $|0\rangle$, while the counter-rotating term $\propto a^{\dagger}_{\nu\ell m\lambda}\sigma_{eg}$ survives. This is the curved-space counterpart of the UDW vacuum excitation via anti-resonant terms, and is the microscopic origin of the emission channel that populates the mode in our master equation. Third and finally, a static detector at fixed $r$ in the Unruh picture samples a local temperature $\propto 1/\sqrt{f(r)}$ Tjoa:2020eqh. Our atoms are not static; they free-fall across the near-horizon region. The observable is the escaping flux at $\mathscr{I}^{+}$, obtained by sewing the near-horizon thermal kernel (which fixes detailed balance) to the global scattering problem (which provides $\Xi_{\ell}(\nu)$ and the threshold). The result is the same thermal ratio (119) multiplied by smooth, channel-dependent prefactors. That is why the entropy law
$\frac{dS_{p}}{dt}=4\pi k_{B}\sum_{\nu,\ell}\nu\,\dot{\bar{n}}_{\nu}(\ell)$ (120)
@@
$\frac{dS_{p}}{dt}=\frac{k_{B}c^{3}}{4\hbar G}\,\dot{A}_{V}$ (121)
hold independent of vector mass and polarization content: the Proca details flow only into $\dot{\bar{n}}_{\nu}$ through thresholds and greybody transmissions Bekenstein:1973ur; Hawking:1975vcx; Bardeen:1973gs.

### v2

@@ VII.2 Area-entropy relation with Proca modifications
We now relate the entropy flux carried by the escaping Proca quanta to the change of the black hole area. Summing (VII.7) over all cavity-admitted channels gives the total entropy rate
$\frac{dS_{p}}{dt}=4\pi k_{B}\sum_{\nu,\ell}\nu\,\dot{\bar{n}}_{\nu}(\ell).$ (VII.8)
@@
$\dot{E}_{V}=\hbar\,\frac{c}{r_{g}}\sum_{\nu,\ell}\nu\,\dot{\bar{n}}_{\nu}(\ell)=\hbar\,\frac{c^{3}}{2GM}\sum_{\nu,\ell}\nu\,\dot{\bar{n}}_{\nu}(\ell).$ (VII.9)
@@
$\dot{A}=\frac{32\pi G^{2}M}{c^{4}}\,\dot{M}.$ (VII.10)
Combining (VII.8)–(VII.10) eliminates the mode sum and yields
$\frac{dS_{p}}{dt}=-\,\frac{k_{B}c^{3}}{4\hbar G}\,\dot{A}.$ (VII.11)
@@
$\frac{dS_{p}}{dt}\;=\;\frac{k_{B}c^{3}}{4\hbar G}\,\dot{A}_{V}\;.$ (VII.13)
Equation (VII.13) is universal: the proportionality constant depends only on $(G,\hbar,c,\,k_{B})$ and the near-horizon mapping that fixed detailed balance, while all Proca-specific physics (mass threshold $\nu\geq m_{V}$, polarization sums, axial/polar greybody factors, and detector realization) enters solely through $\dot{A}_{V}$ via the net flux $\sum_{\nu,\ell}\nu\,\dot{\bar{n}}_{\nu}(\ell)$ in (VII.9). In the massless limit $m_{V}\to 0$ the same relation holds, with $\dot{A}_{V}$ computed from the corresponding threshold-free spectrum.
@@ VII.3 Spectral signatures and limits
We distill the qualitative and asymptotic features of the Proca spectrum that enter the entropy flux through $\dot{\bar{n}}_{\nu}$ and, hence, $\dot{S}_{p}$ in (VII.7)–(VII.13). The thermal denominator is general; all structure resides in the smooth prefactor and the kinematic threshold $\nu\geq m_{V}$. In particular, the suppression near $\nu\simeq m_{V}$ is a threshold/scattering effect controlled by $k_{\infty}(\nu)$ and $\Xi_{\ell}(\nu)$; it is logically distinct from the large-$\omega$ asymptotic expansion used to derive the universal near-horizon kernel.
@@
The massless limit $m_{V}\to 0$ removes the threshold and sends $k_{\infty}\to\nu$, returning the scalarlike normalization $1/\nu$ (up to polarization weights). The dipole model passes smoothly to the standard electromagnetic response. These observations also clarify the semirealistic physical meaning of the Proca extension. The new ingredients are not limited to extra algebraic structure: the massive vector field predicts a hard low-frequency gap, a polarization/channel hierarchy close to threshold, and coupling-dependent longitudinal/transverse weights, all of which are absent in the scalar HBAR problem. In a cavity-resolved or detector-based realization, these quantities would be the direct observables distinguishing massive spin-1 acceleration radiation from its scalar or Maxwell counterparts. The present analysis therefore identifies which parts of the signal are universal consequences of near-horizon kinematics and which parts encode the genuinely new physics of a Proca field. The current model, while formulated naturally for Proca, still yields a smooth limit at the level of observable probabilities and rates. Finally, all results here rest on the large-$\omega$ expansion that produced the universal near-horizon integral; corrections scale as $\mathcal{O}(\nu/\omega)$, reshaping only the smooth prefactor and leaving intact the Planckian denominator and detailed balance. Consequently, the entropy flux $\dot{S}_{p}=4\pi k_{B}\sum_{\nu,\ell}\nu\dot{\bar{n}}_{\nu}(\ell)$ inherits the same signatures: delayed onset at $\nu\approx m_{V}$, gentle turn-on dominated by polar $\ell=0$, and a universal thermal tail at higher frequencies.
These signatures (sharp threshold, channel-dependent turn-on, polarization-weighted prefactors, and a universal thermal tail) are the distinctive Proca imprints in the HBAR entropy law (VII.13).
@@ VIII.1 Large-$\omega$ expansion and validity
$\mathcal{A}_{\bullet}(\nu,\ell,m,\lambda)\;=\;\int d\tau\;\mathcal{G}_{\bullet}(\tau,\nu,\ell,m,\lambda)\,e^{\,i\Phi(\tau,\nu)}.$ (VIII.1)
@@
$\Phi(\tau,\nu)=\omega\tau+\nu\,\left[t(\tau)-r_{*}(\tau)\right]+\Phi_{0}(\nu),$ (VIII.2)
@@
$\mathcal{I}(\nu,\omega)\;\equiv\;\int_{0}^{\infty}dy\,y^{-2i\nu}\,e^{-i\omega y}\;=\;(i\omega)^{-(1-2i\nu)}\Gamma(1-2i\nu),$ (VIII.3)
@@
$|\mathcal{I}(\nu,\omega)|^{2}=\frac{4\pi\,\nu}{\omega^{2}}\,\frac{1}{e^{4\pi\nu}-1}.$ (VIII.4)
The large-$\omega$ control arises because the integrand oscillates rapidly on a scale $y\sim\omega^{-1}$. Provided the envelope is smooth on this scale, we may expand both the smooth prefactor and the regular part of the phase about the horizon. Writing
$\mathcal{G}_{\bullet}(\tau_{h}-y,\nu,\cdot)\,e^{\,i\nu\mathcal{O}(y)}=\sum_{n=0}^{\infty}c_{n}^{(\bullet)}(\nu,\cdot)\,y^{n},$ (VIII.5)
the amplitude becomes
$\mathcal{A}_{\bullet}$ $=\sum_{n=0}^{\infty}c_{n}^{(\bullet)}(\nu,\cdot)\,\mathcal{I}_{n}(\nu,\omega),$
$\mathcal{I}_{n}(\nu,\omega)$ $\equiv\int_{0}^{\infty}dy\,y^{n-2i\nu}e^{-i\omega y}.$ (VIII.6)
These moments are evaluated exactly as
$\mathcal{I}_{n}(\nu,\omega)=\frac{\Gamma(n+1-2i\nu)}{(i\omega)^{\,n+1-2i\nu}}=\mathcal{I}(\nu,\omega)\,\frac{(1-2i\nu)_{n}}{(i\omega)^{n}},$ (VIII.7)
where $(a)_{n}$ is the Pochhammer symbol. Therefore
$\mathcal{A}_{\bullet}=\mathcal{I}(\nu,\omega)\left[c_{0}^{(\bullet)}+\frac{(1-2i\nu)c_{1}^{(\bullet)}}{i\omega}+\mathcal{O}\!\left(\frac{1+\nu^{2}}{\omega^{2}}\right)\right],$ (VIII.8)
and, correspondingly,
$P_{\bullet}(\nu,\ell)=P_{\bullet}^{(0)}(\nu,\ell)\left[1+\mathcal{O}\!\left(\frac{1+|\nu|}{\omega}\right)\right].$ (VIII.9)
Thus the asymptotic parameter is $1/\omega$, while the emitted-mode frequency $\nu$ enters only through smooth polynomial coefficients multiplying the same universal kernel $\mathcal{I}(\nu,\omega)$.
This makes explicit that the low-frequency suppression discussed in Sec. VII.3 is not produced by a breakdown of the large-$\omega$ expansion. Rather, it comes from the escaping-mode prefactor
$P_{\mathrm{exc,esc}}^{\bullet}(\nu,\ell)\propto\frac{\Theta(\nu-m_{V})}{k_{\infty}(\nu)}\,\Xi_{\ell}(\nu)\,\frac{1}{e^{4\pi\nu}-1},$ (VIII.10)
with $k_{\infty}(\nu)=\sqrt{\nu^{2}-m_{V}^{2}}$ and $\Xi_{\ell}(\nu)\to 0$ as $\nu\to m_{V}^{+}$. In particular, if $\Xi_{\ell}(\nu)\sim k_{\infty}^{2\ell+1}$ near threshold, then $P_{\mathrm{exc,esc}}^{\bullet}(\nu,\ell)\propto k_{\infty}^{2\ell}$, so the turn-on at low $\nu$ is a threshold/greybody effect rather than a failure of the $1/\omega$ expansion.
The universal Planckian denominator and the detailed-balance ratio originate from the common factor $\mathcal{I}(\nu,\omega)$, i.e. from the logarithmic near-horizon mapping of $t-r_{*}$. Subleading terms only renormalize the smooth channel-dependent prefactor and therefore affect absolute rates and relaxation times, not the leading geometric thermal kernel. The regime of validity is
$\omega\,\Delta\tau_{\rm cav}\gg 1,\qquad\Delta\nu\ll\min\{\nu,\omega\},\qquad\frac{1+|\nu|}{\omega}\ll 1.$ (VIII.11)
@@
$\omega\,\delta\tau_{\rm cav}\gg 1\quad\Rightarrow\quad\text{single-pass, near-horizon dominance.}$ (VIII.12)
@@ VIII.2 Gauge and constraints
The Proca theory is massive and therefore has no gauge symmetry. The covariant Lorenz condition $\nabla_{\mu}A^{\mu}=0$ follows dynamically from the equations of motion, leaving three propagating degrees of freedom. This structure percolates cleanly through mode expansion and detector couplings. Asymptotically, the three polarizations satisfy $k\!\cdot\!\varepsilon^{(\lambda)}=0$, and the projector is [50]
$\sum_{\lambda=1}^{3}\varepsilon^{(\lambda)}_{\mu}\varepsilon^{(\lambda)}_{\nu}=-\eta_{\mu\nu}+\frac{k_{\mu}k_{\nu}}{m_{V}^{2}}.$ (VIII.13)
Near the horizon, the axial/polar split maps to three independent master variables with plane wave behavior $e^{-i\nu(t-r_{*})}$; the constraint does not modify the universal phase.
@@
$u\!\cdot\!u+\frac{(u\!\cdot\!k)^{2}}{m_{V}^{2}}=\,-1+\frac{\Omega^{2}}{m_{V}^{2}},$ (VIII.14)
@@
$E_{\mu}\equiv F_{\mu\nu}u^{\nu}=i\left[k_{\mu}(A\!\cdot\!u)-(k\!\cdot\!u)\,A_{\mu}\right]$ (VIII.15)
automatically satisfies $u^{\mu}E_{\mu}=0$. Both pieces contribute at the amplitude level; after summing polarizations the weight is built from contractions with the projector in (VIII.13). Orientation averaging of $d^{\mu}$ in the atom’s rest frame produces a smooth, $\mathcal{O}(1)$ function of $\nu$.
Finally, for the massless limit, as $m_{V}\to 0$, the projector (VIII.13) becomes singular, but all observables remain finite: the combination of unit-flux normalization, $\Xi_{\ell}(\nu)$, and the longitudinal/transverse balance yields a smooth limit in probabilities and rates. Physically, the dipole model tends to the familiar electromagnetic response; the current model, while not gauge-invariant in the $m_{V}=0$ theory, approaches the same observables because the longitudinal contribution decouples as $\nu$ grows and $\Xi_{\ell}\to 1$.
@@ VIII.3 Comparison with Unruh/DeWitt logic
Our calculation is the black hole (Killing/Rindler) analogue of the Unruh-DeWitt (UDW) detector response. Three points make the connection precise. First, in the UDW setting, a uniformly accelerated detector with proper acceleration $a$ sees a thermal spectrum with temperature $T=a/(2\pi)$, traced to the KMS periodicity of the Wightman function in imaginary proper time. Here, along the free fall trajectory, the near-horizon mapping
$t(\tau)-r_{*}(\tau)=-\,2\ln(\tau_{h}-\tau)+C_{h}+\cdots$ (VIII.16)
acts as a local Rindler time. The logarithm produces a branch cut whose discontinuity across imaginary $2\pi$ shifts enforces [51, 27]
$\frac{P_{\rm exc}(\nu)}{P_{\rm abs}(\nu)}=e^{-4\pi\nu},$ (VIII.17)
exactly the KMS ratio with surface gravity $\kappa=\frac{1}{2}$ in our dimensionless units. This universality explains why spin, mass, and detector realization cannot change the thermal denominator. Second, in the Boulware vacuum, the rotating term $\propto a_{\nu\ell m\lambda}\sigma_{eg}$ vanishes on $|0\rangle$, while the counterrotating term $\propto a^{\dagger}_{\nu\ell m\lambda}\sigma_{eg}$ survives. This is the curved-space counterpart of the UDW vacuum excitation via antiresonant terms, and is the microscopic origin of the emission channel that populates the mode in our master equation. Third and finally, a static detector at fixed $r$ in the Unruh picture samples a local temperature $\propto 1/\sqrt{f(r)}$ [49]. Our atoms are not static; they free fall across the near-horizon region. The observable is the escaping flux at $\mathscr{I}^{+}$, obtained by sewing the near-horizon thermal kernel (which fixes detailed balance) to the global scattering problem [which provides $\Xi_{\ell}(\nu)$ and the threshold]. The result is the same thermal ratio (VIII.17) multiplied by smooth, channel-dependent prefactors. That is why the entropy law
$\frac{dS_{p}}{dt}=4\pi k_{B}\sum_{\nu,\ell}\nu\,\dot{\bar{n}}_{\nu}(\ell)$ (VIII.18)
@@
$\frac{dS_{p}}{dt}=\frac{k_{B}c^{3}}{4\hbar G}\,\dot{A}_{V}$ (VIII.19)
hold independent of vector mass and polarization content: the Proca details flow only into $\dot{\bar{n}}_{\nu}$ through thresholds and greybody transmissions [7, 29, 5].